OpenStudy (anonymous):

Solve for the unknown: G – 24 = 75

OpenStudy (anonymous):

75+24=?

OpenStudy (anonymous):

g=99

OpenStudy (anonymous):

^Correct

OpenStudy (anonymous):

we cant give you an exact answer but you will have to solve that for the equasion @JoStudy gave you

OpenStudy (anonymous):

g-24=75 +24+24 = g 99